### v1

# The polynomial set associated with a fixed number of matrix-matrix multiplications

Abstract
We consider the problem of computing matrix polynomials $p(X)$, where $X$ is a large matrix, with as few matrix-matrix multiplications as possible. More precisely, let $\Pi_{2^{m}}^{*}$ represent the set of polynomials computable with $m$ matrix-matrix multiplications, but with an arbitrary number of matrix additions and scaling operations. We characterize this set through a tabular parameterization. By deriving equivalence transformations of the tabular representation, we establish new methods that can be used to construct elements of $\Pi_{2^{m}}^{*}$ and determine general properties of the set. The transformations allow us to eliminate variables and prove that the dimension is bounded by $m^{2}$. Numerical simulations suggest that this is a sharp bound. Consequently, we have identified a parameterization, which, to our knowledge, is the first minimal parameterization. Furthermore, we conduct a study using computational tools from algebraic geometry to determine the largest degree $d$ such that all polynomials of that degree belong to $\Pi_{2^{m}}^{*}$, or its closure. In many cases, the computational setup is constructive in the sense that it can also be used to determine a specific evaluation scheme for a given polynomial.
1 Introduction
The computation of matrix functions, in the sense of Higham [13], is a classical problem in numerical linear algebra. We define a matrix function as an extension of a scalar function from $f:\mathbb{C}\rightarrow\mathbb{C}$ to matrices, i.e., $f:\mathbb{C}^{n\times n}\rightarrow\mathbb{C}^{n\times n}$. Important matrix functions like $e^{X}$, $\text{sign}(X)$, and $\sqrt{X}$ are crucial in various fields such as linear ODEs [18], control theory [5], network analysis [7], and quantum chemistry [22]; see [13, Chapter 2] for more applications. Further applications relevant for our setting, appear in problems where the action of a matrix function on a vector $b\in\mathbb{C}^{n}$, i.e., $f(X)b$ is needed, see [9].
In this paper we study methods for computing matrix functions when the input matrix $X\in\mathbb{C}^{n\times n}$ is a very large and dense matrix. In particular, we consider a family of methods that only utilize the following operation types:
  • Linear combination: $Z\leftarrow\alpha X+\beta Y$
  • Multiplication: $Z\leftarrow X\cdot Y$.
A direct consequence of only considering these basic operation types is that this family of methods compute matrix polynomials. Another direct consequence is that the first operation type can be viewed as free in terms of computational cost, since the computational complexity is $O(n^{2})$ and $O(n^{3})$ respectively for the two considered operation types.
We want to study what polynomials can be computed with a given cost. Since, the cost is essentially given by the number of matrix-matrix multiplications, methods with a given fixed cost corresponding to $m$ multiplications, form evaluations of polynomials in the set
[…]
where $\Pi_{d}$ is the set of all polynomial of degree $d$ in the usual sense. Here, $\widebar{\Pi}_{d}$ is the closure of $\Pi_{d}$ which is the set of all polynomial of grade $d$, that is, polynomials of degree $\leq d$. In our analysis, we will need to be extra careful concerning the distinction between $\Pi_{k}$ and $\widebar{\Pi}_{k}$. In fact, $\Pi_{2^{m}}^{*}$ is a semi-algebraic set, as we concretize in Section 2.
In an application involving a function $f$, which is not necessarily a polynomial, we often want compute an approximation $p$ of $f$ using a fixed cost. Hence, we want to find the best polynomial approximation in $\Pi_{2^{m}}^{*}$. This is in complete contrast to the classical problem of (scalar) approximation theory where we want to find the best approximation of a given fixed degree. Since $\Pi_{2^{m}}^{*}\neq\widebar{\Pi}_{2^{m}}$, the fixed degree and fixed cost approximation problems are different in character. In order to construct methods for the fixed cost approximation problem we need to understand the set $\Pi_{2^{m}}^{*}$. The objective of this paper is to characterize this set.
Techniques to keep the number of matrix-matrix multiplications low has been studied for decades in the context of matrix functions and matrix polynomials. In Paterson and Stockmeyer’s seminal paper [19] they present an algorithm to compute $p(X)$ for $p\in\widebar{\Pi}_{d}$ in $m=\mathcal{O}(\sqrt{d}$) matrix-matrix multiplications, which was made more precise in [8]. Since they also show that dim$(\Pi_{2^{m}}^{*})=\mathcal{O}(m^{2})$ = $\mathcal{O}(d)$, Paterson and Stockmeyer’s algorithm is optimal in the ordo sense.
The fact that the Paterson and Stockmeyer algorithm is improvable, in particular for moderate $m$, has served as the motivation of a number of recent works. For example, Sastre [24] showed how to compute most polynomials of degree $8$ using only three multiplications, and most polynomials of degree $12$ using only four multiplications (in contrast to the Paterson-Stockmeyer algorithm that handle degrees $6$ and $9$ respectively). The work has been extended in various ways, for example [25], illustrating how an approximation of the Taylor expansion of the exponential can be computed with $m=4$ multiplications; called the approximation 15+ corresponding to a polynomial of degree 16 matching all Taylor coefficients except the last, i.e., the coefficient for $X^{16}$. Corresponding ideas for rational functions were explored in [23]. Further specialziation of the algorithms, in particular for scaling-and-squaring [17, Section 7], has been investigated for the matrix exponential [26], [27], as well as the matrix cosine [28]. Using ideas of efficient polynomial evaluation combined with rational approximations have been used [2, 3], with particular focus on preservation of Lie algebra properties for the matrix exponential. In general, such efficient polynomial methods are (at least in theory) better than the standard implementation of the matrix exponential using a Padé approximant combined with the scaling-and-squaring algorithm [1], although further research is needed to conclusively support this claim.
In the terminology of our framework, considerable parts of the results in the literature discussed above, are examples involving polynomials that are elements of $\Pi_{2^{m}}^{*}$. Despite this research attention, the understanding of this set is far from complete. The basis of our analysis is an evaluation scheme (further explained in Section 2) that gives a parameterization of $\Pi_{2^{m}}^{*}$ with parameters given by a triplet $(A,B,c)$, where $A$ and $B$ are matrices and $c$ a vector. One triplet $(A,B,c)$ corresponds to one polynomial $p\in\Pi_{2^{m}}^{*}$. There are several ways to obtain the same polynomial, i.e., a different triplet $(\hat{A},\hat{B},\hat{c})$ may lead to the same polynomial $p$.
The first set of new results are equivalence transformations. In Section 3, we present procedures to transform a triplet $(A,B,c)$ into another triplet $(\hat{A},\hat{B},\hat{c})$ that yield the same polynomial $p\in\Pi_{2^{m}}^{*}$. Several conclusions can be drawn from the transformations. For example, we prove that the first column of the matrices $A$ and $B$ can be selected as zero without loss of generality. More complex transformations reveal that the $(2,2)$ element of $A$ can also be set to zero. Moreover, we establish a transformation related to the third multiplication, which includes an algebraic condition on the elements of $A$ and $B$. Further analysis shows that the $(3,3)$ element of $A$ can be coupled in simple way to the $(3,3)$ element of $B$ without loss of generality.
Further conclusions from the equivalence transformations are presented in Section 4. Using the transformations we can reduce the number of free parameters that parameterize $\Pi_{2^{m}}^{*}$ and conclude that
$\dim(\Pi_{2^{m}}^{*})\leq m^{2}\textrm{ for }m\geq 3.$
This is sharper than bounds presented in, e.g., [19]. Based on a study of the numerical Jacobian, we conjecture that this bound cannot be improved, suggesting that our work present the (to our knowledge) first minimal parameterization of $\Pi_{2^{m}}^{*}$.
Further results are presented (in Section 5) regarding the problem of finding included polynomial subsets, specifically determining
(2) $\max\{d:\Pi_{d}\subset\widebar{\Pi_{2^{m}}^{*}}\}.$
Using the transformations and tools from computational algebraic geometry, we solve this problem for $m=4$ and provide conjectures supported by strongly indicative computational results in high precision arithmetic for $m=5$, $m=6$, and $m=7$. Our solutions to (2) are $12$, $20$, $32$, $30$, and $42$ for $m=4$, $m=5$, $m=6$ (complex arithmetic), $m=6$ (real arithmetic), and $m=7$ (complex arithmetic), respectively. The numerical experiments are reproducible and provided in a publicly available repository, including generated code for matlab and Julia.
Although we focus on the determination of dimension and the solution to (2), several of the results seem valuable beyond these specific research questions. The transformation by themselves are constructive, and could be used to improve other properties of the evaluation scheme, e.g., numerical stability. The study of the polynomial subset in Section 5 is constructive. We provide evaluation schemes for the truncated Taylor expansion of the matrix exponential for the given degrees. For example, we compute the degree-$30$ expansion in $m=6$ multiplications, while previous works appear to require $7$ multiplications. Moreover, the software that we provide is able to compute the evaluation scheme coefficients for many given polynomials.
2 Evaluation scheme
2.1 Definition of the evaluation scheme
Fundamental to the results of this paper is an evaluation scheme to construct elements of $\Pi_{2^{m}}^{*}$. Such evaluation schemes have been presented in various forms in [19, Section 2], [2, section 3], and also [14], where in the latter they are referred to as degree-optimal polynomials. The scheme involves parameters stored in matrices $A$ and $B$, and vector $c$. Matrices $A$ and $B$ contain coefficients for linear combinations, where each row $k$ corresponds to the linear combinations associated with the $k$th multiplication. As we perform $m$ multiplications, these matrices have $m$ rows. The vector $c$ contains coefficients for linear combinations that are used after the multiplications have been completed.
In particular, consider the triplet $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ and associate the following sequence of operations involving precisely $m$ matrix-matrix multiplications. Define $Q_{1}=I$ and $Q_{2}=X$. Then, iterate the process to generate $Q_{3},\ldots,Q_{m+2}$ using the elements of matrices $A$ and $B$ as follows:
(3) $Q_{3}$ $=(a_{11}Q_{1}+a_{12}Q_{2})(b_{11}Q_{1}+b_{12}Q_{2})$
$Q_{4}$ $=(a_{21}Q_{1}+a_{22}Q_{2}+a_{23}Q_{3})(b_{21}Q_{1}+b_{22}Q_{2}+b_{23}Q_{3})$
$Q_{5}$ $=(a_{31}Q_{1}+a_{32}Q_{2}+a_{33}Q_{3}+a_{34}Q_{4})(b_{31}Q_{1}+b_{32}Q_{2}+b_{33}Q_{3}+b_{34}Q_{4})$
$\vdots$
$Q_{m+2}$ $=(a_{m,1}Q_{1}+\cdots+a_{m,m+1}Q_{m+1})(b_{m,1}Q_{1}+\cdots+b_{m,m+1}Q_{m+1}).$
Furthermore, compute the output of the scheme by forming the linear combination corresponding to the $c$-vector:
(4) $p(X)=c_{1}Q_{1}+c_{2}Q_{2}+\cdots+c_{m+2}Q_{m+2}.$
Thus, a given triple $(A,B,c)$, with $A$ and $B$ being Hessenberg matrices, defines a polynomial $p\in\Pi_{2^{m}}^{*}$.
Note that all evaluation schemes can be expressed in this manner, given that for each multiplication step, we use all preceding information accessible at no additional computational expense. For instance, in the second multiplication step, we determine $Q_{4}$ by combining linearly the prior multiplications $Q_{3}$ along with the inputs $Q_{2}$ and $Q_{1}$. Hence, a triplet $(A,B,c)$ parameterizes $\Pi_{2^{m}}^{*}$ through the expressions in (3) and (4). Detailed examples illustrating the representation of monomial evaluation or Horner evaluation can be found in [14].
Examples of instances of $\Pi_{2^{m}}^{*}$
To see how the tables represent standard evaluation schemes, such as monomial evaulation, Horner evaluation, and Paterson-Stockmeyer evaluation, see [14, Section 3]. In the following, we show how to evaluate $X^{7}+\epsilon X^{8}$ in three multiplications. Consider the coefficient triplet:
(8) $[A|B]$ $=\left[\begin{array}[]{c c c c | c c c c}0&1&&&0&1&&\\ 0&\frac{1}{2\epsilon}&1&&0&0&1&\\ 0&\frac{7+8\epsilon^{2}-16\epsilon^{4}}{128\epsilon^{5}}&-\frac{1}{8\epsilon^{2}}-\frac{1}{2}&1&0&\frac{-7+8\epsilon^{2}+16\epsilon^{4}}{128\epsilon^{5}}&-\frac{1}{8\epsilon^{2}}+\frac{1}{2}&1\end{array}\right]$
(10) $c$ $=\left[\begin{array}[]{ccccc}0&0&\frac{49-288\epsilon^{4}+256\epsilon^{8}}{16384\epsilon^{9}}&\frac{-5+16\epsilon^{4}}{64\epsilon^{3}}&\epsilon\end{array}\right].$
For this triplet $(A,B,c)$, we obtain the polynomial $p(X)=X^{7}+\epsilon X^{8}$ using the evaluations specified in (3)-(4). We observe that $X^{7}$ is as a limit point of $\Pi_{2^{3}}^{*}$ corresponding to $\epsilon\rightarrow 0$, although $X^{7}$ requires four multiplications. This example illustrates the complexity of $\Pi_{2^{m}}^{*}$ for $m=3$. For higher $m$, this complexity becomes even more intricate.
2.2 The semi-algebraic set perspective of $\Pi_{2^{m}}^{*}$
For our study, we need concepts from algebraic geometry. Let $\mathbb{K}$ be the field of the parameters in the evaluation, i.e., in most applications $\mathbb{K}=\mathbb{C}$ or $\mathbb{K}=\mathbb{R}$. Let $\mathbb{X}$ be the product space associated with the parameters. That is $\mathbb{X}:=\mathbb{K}^{m(m+3)/2}\times\mathbb{K}^{m(m+3)/2}\times\mathbb{K}^{m+2}=\mathbb{K}^{s}$, where $s=m^{2}+4m+2$ is the number of parameters in the evaluation.
If we denote the evaluation scheme (3)-(4) by the map $\Phi:(A,B,c)\mapsto p\in\widebar{\Pi}_{d}$, we can describe $\Pi_{2^{m}}^{*}$ as its image
(11) $\Phi(\mathbb{X})=\Pi_{2^{m}}^{*}.$
By construction, $\Phi$ depends polynomially on the parameters. This can be concluded from the property that the image of a smooth map, with a vector space as its domain, is inherently semi-algebraic.
The set is indeed semi-algebraic and not necessarily an algebraic variety. Let $\widebar{\cdot}$ denote topological closure in a Zariski sense. Then, in relation to the example (8) above, we see that $\epsilon=0$ is not a valid choice, and indeed $X^{7}\not\in\Pi_{2^{3}}^{*}$. However, since $\epsilon=0$ is a limit point, we have $X^{7}\in\widebar{\Pi_{2^{3}}^{*}}=\widebar{\Pi}_{2^{3}}$; at least for $\mathbb{K}=\mathbb{C}$ due to [32, Corollary 4.20].
If $\varphi_{0},\ldots,\varphi_{d}$ denote a basis of the ambient space $\mathbb{K}[x]_{d}$, e.g., as in our simulations a monomial basis, and let $J$ be the Jacobian of the map $\Phi$ with an output expressed in this basis. Following the standard definitions in algebraic geometry [29], the dimension of a semi-algebraic set is given by the rank of $J$ at any non-singular point. Most points in the image of a smooth map are non-singular; therefore for essentially all $(A,B,c)$ we have
[…]
where $J\in\mathbb{K}^{(d+1)\times s}$ and $d=2^{m}$ and $s$ is the number of parameters in the method.
The dimension of $\Pi_{2^{m}}^{*}$ is always bounded by the number of free parameters, and in this setting it can be concluded from the size of $J$ that
(12) $\dim\Pi_{2^{m}}^{*}\leq s=m^{2}+4m+2$
In this paper we improve this bound by reducing the number of free parameters and in Section 4 conclude that $m^{2}$ seems to be the exact dimension, for $m>3$.
Here we note that the dimension may be different for $\mathbb{K}=\mathbb{C}$ or $\mathbb{K}=\mathbb{R}$. Most of the results in this paper are valid for either choice, but in situations where we explicitly refer to the Zariski closure, we assume $\mathbb{K}=\mathbb{C}$ unless otherwise stated. In the simulations in Section 5 we focus on $\mathbb{K}=\mathbb{C}$ with some additional observations for $\mathbb{K}=\mathbb{R}$.
3 Equivalence transformations
3.1 Substitution transformations
The following theorems describe transformations of the evaluation scheme $(A,B,c)$ in a way that the output polynomial $p(X)$ is unchanged. The modified evaluation scheme will be denote $(\hat{A},\hat{B},\hat{c})$ and the corresponding $Q$-matrices will be denoted $\hat{Q}_{1},\ldots,\hat{Q}_{m+1}$. The first transformation follows from a change of variables, where we rescale one of the $Q$-coefficients. For example if we set
$\hat{Q}_{4}=\alpha Q_{4}$
the tables (3) are changed by
$\hat{Q}_{4}=(\alpha a_{21}Q_{1}+\alpha a_{22}Q_{2}+\alpha a_{23}Q_{3})(b_{21}Q_{1}+b_{22}Q_{2}+b_{23}Q_{3}).$
In order to keep $Q_{5},\ldots,Q_{m+2}$ unchanged, we can cancel the transformation by scaling the coefficient in the corresponding term, e.g.,
$Q_{5}=(a_{31}Q_{1}+a_{32}Q_{2}+a_{33}Q_{3}+\alpha^{-1}a_{34}\hat{Q}_{4})(b_{31}Q_{1}+b_{32}Q_{2}+b_{33}Q_{3}+\alpha^{-1}b_{34}\hat{Q}_{4}).$
This implies that without changing the polynomial $p$, we may rescale one row in $A$ and one column in $A$ and $B$. This is formalized in the following theorem, where we omit index on elements of $A$ and $B$ that are unchanged in $\hat{A}$ and $\hat{B}$. From symmetry, we note that the same result holds for $A$ and $B$ switched.
Theorem 3.1.
Let $p$ be the polynomial associated with $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ and $\hat{p}$ be the polynomial associated with $(\hat{A},\hat{B},\hat{c})$, where
(13) $\hat{A}$ $=\left[\begin{array}[]{c c c c c c c c c c}a&a&\\ \vdots&&\ddots\\ a&\cdots&\cdots&a\\ \hat{a}_{k,1}&\cdots&\cdots&\cdots&\hat{a}_{k,k+1}\\ a&\cdots&\cdots&\cdots&a&\hat{a}_{k+1,k+2}\\ \vdots&&&&\vdots&\vdots&a\\ \vdots&&&&\vdots&\vdots&\vdots&\ddots\\ a&\cdots&\cdots&\cdots&a&\hat{a}_{m,k+2}&a&\cdots&a\\ \end{array}\right]$
$\hat{B}$ $=\left[\begin{array}[]{c c c c c c c c c c}b&b&\\ \vdots&&\ddots\\ \phantom{x}b\phantom{x}&\cdots&\cdots&b\\ b&\cdots&\cdots&\cdots&\phantom{xx}b\phantom{xx}\\ b&\cdots&\cdots&\cdots&b&\hat{b}_{k+1,k+2}\\ \vdots&&&&\vdots&\vdots&b\\ \vdots&&&&\vdots&\vdots&\vdots&\ddots\\ b&\cdots&\cdots&\cdots&b&\hat{b}_{m,k+2}&b&\cdots&b\\ \end{array}\right]$
$\hat{c}$ $=\;\left[\begin{array}[]{c c c c c c c c c}\phantom{x}c_{1}&\cdots&\cdots&\cdots&c_{k+1}&\alpha^{-1}c_{k+2}&c_{k+3}&\cdots&c_{m+2}\end{array}\right],$
and
(14a) $\hat{a}_{k,j}$ $=$ $\alpha a_{k,j},\;\;\;\qquad j=1,\ldots,k+1$
(14b) $\hat{a}_{i,k+2}$ $=$ $\alpha^{-1}a_{i,k+2},\;\;\;i=k+1,\ldots,m$
(14c) $\hat{b}_{i,k+2}$ $=$ $\alpha^{-1}b_{i,k+2},\;\;\;i=k+1,\ldots,m.$
Then,
(15) $\hat{p}=p,$
for any $k\in\left[1,\ldots,m\right]$ and $\0\alpha\in\mathbb{C}\backslash\{0\}$.
Proof 3.2.
By definition, we have $\hat{Q}_{1}=Q_{1}=I$ and $\hat{Q}_{2}=Q_{2}=X$. The proof is based on establishing the following three statements:
(16a) $\hat{Q}_{i+2}$ $=$ $Q_{i+2}\qquad\text{for }i=1,\ldots,k-1,$
(16b) $\hat{Q}_{k+2}$ $=$ $\alpha Q_{k+2}$
(16c) $\hat{Q}_{i+2}$ $=$ $Q_{i+2}\qquad\text{for }i=k+1,\ldots,m+2.$
Together with the definition of $\hat{c}$ in (13), this implies that the conclusion of the theorem, stated in (15), holds.
To prove (16a), we observe that since the first $k-1$ rows of $\hat{A}$ and $\hat{B}$ are unchanged, the variables $\hat{Q}_{1},\ldots,\hat{Q}_{k+1}$ are also unchanged.
To prove (16b), we consider row $k$, i.e., the first row where there are changes. Inserting (14a) in the definition of $\hat{Q}_{k+1}$ yields a scaling,
$\hat{Q}_{k+2}$ $=(\alpha a_{k,1}Q_{1}+\cdots+\alpha$ $a_{k,k+1}Q_{k+1}$ $)(b_{k,1}Q_{1}+\cdots+b_{k,k+1}Q_{k+1}$ $)=\alpha Q_{k+2}$
which proves (16b).
To prove (16c), we first consider (16c) for $i=k+1$. Inserting (14b) in the definition of $\hat{Q}_{k+3}$ results in a cancellation
$\hat{Q}_{k+3}$ $=(a_{k+1,1}Q_{1}+\ldots+\alpha^{-1}a_{k+1,k+2}\hat{Q}_{k+2}$ $)(b_{k+1,1}Q_{1}+\ldots+\alpha^{-1}b_{k+1,k+2}\hat{Q}_{k+2}$ )))
$=Q_{k+3},$
where the last equality follows from (16b). The general statement follows from induction.
In a similar fashion we carry out a change of variables where we add one given value $\alpha$ to one of the elements in the first column. For example, adding $\alpha$ to the coefficient corresponding to $Q_{1}=I$, in the first multiplication, leads to
$\hat{Q}_{3}=\left((a_{11}+\alpha)Q_{1}+a_{12}Q_{2}\right)\left(b_{11}Q_{1}+b_{12}Q_{2}\right)$
which can be expanded as
(17) $\hat{Q}_{3}:=Q_{3}+\alpha(b_{11}Q_{1}+b_{12}Q_{2}).$
Let $A_{k}$ and $B_{k}$ be the factors that form $Q_{k}$, i.e., $Q_{k}=A_{k}B_{k}$. In order to keep $Q_{4},\ldots,Q_{m+2}$ unchanged, we keep both factors in $\hat{Q}_{4}=Q_{4}=A_{4}B_{4}$ unchanged by compensating for the transformation (17):
(18) $A_{4}$ $=$ $(a_{21}-\alpha a_{23}b_{11})Q_{1}+(a_{22}-\alpha a_{23}b_{12})Q_{2}+a_{23}\hat{Q_{3}}$
(19) $B_{4}$ $=$ $(b_{21}-\alpha b_{23}b_{11})Q_{1}+(b_{22}-\alpha b_{23}b_{12})Q_{2}+b_{23}\hat{Q_{3}}.$
Hence, a modification in the coefficient corresponding to $Q_{1}$ can be compensated by modifying coefficients in rows below the modification. This can be applied to an arbitrary row $k$ and arbitrary $\alpha$.
Theorem 3.3.
Let $p$ be the polynomial associated with $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ and $\hat{p}$ be the polynomial associated with $(\hat{A},\hat{B},\hat{c})$ where
(20) $\hat{A}$ $=\left[\begin{array}[]{c c c c c c c c c c}a&a&\\ \vdots&&\ddots\\ a&a&\cdots&a\\ a+\alpha&a&\cdots&\cdots&a\\ \hat{a}_{k+1,1}&\hat{a}_{k+1,2}&\cdots&\cdots&\hat{a}_{k+1,k+1}&a\\ \vdots&\vdots&&&\vdots&\vdots&\ddots\\ \hat{a}_{m,1}&\hat{a}_{m,2}&\cdots&\cdots&\hat{a}_{m,k+1}&a&\cdots&a\\ \end{array}\right]$
$\hat{B}$ $=\left[\begin{array}[]{c c c c c c c c c c}b&b&\\ \vdots&&\ddots\\ b&\cdots&\cdots&b\\ \hat{b}_{k+1,1}&\cdots&\cdots&\hat{b}_{k+1,k+1}&b\\ \vdots&&&\vdots&\vdots&\ddots\\ \hat{b}_{m,1}&\cdots&\cdots&\hat{b}_{m,k+1}&b&\cdots&b\\ \end{array}\right]$
$\hat{c}$ $=\;\left[\;\begin{array}[]{c c c c c c c c}\phantom{x}\hat{c}_{1}\phantom{x}&\cdots&\cdots&\phantom{x}\hat{c}_{k+1}\phantom{x}&c&\cdots&\cdots&c\end{array}\right].$
and
(21a) $\hat{a}_{i,j}$ $=$ $a_{i,j}-\alpha a_{i,k+2}b_{k,j},\;\quad j=1,\ldots,k+1,\quad i=k+1,\ldots,m,$
(21b) $\hat{b}_{i,j}$ $=$ $b_{i,j}-\alpha b_{i,k+2}b_{k,j},\;\;\quad j=1,\ldots,k+1,\quad i=k+1,\ldots,m,$
(21c) $\hat{c}_{j}$ $=$ $c_{j}-\alpha c_{k+2}b_{k,j},\;\;\qquad j=1,\ldots,k+1.$
Then,
(22) $\hat{p}=p,$
for any $k\in\left[1,\ldots,m\right]$ and $\alpha\in\mathbb{C}$.
Proof 3.4.
Similarly to the proof of Theorem 3.1, we establish the following three statements:
(23a) $\hat{Q}_{i+2}$ $=$ $Q_{i+2}\quad\text{for }i=1,\ldots,k-1,$
(23b) $\hat{Q}_{k+2}$ $=$ $Q_{k+2}+\alpha B_{k+2}$
(23c) $\hat{Q}_{i}$ $=$ $Q_{i}\quad\text{for }i=k+3,\ldots,m+2.$
Together with the definition of $\hat{c}$ in (21c), this implies that the conclusion of the theorem in equation (22) holds. Relation (23a) follows analogously to the proof in Theorem 3.1.
To prove (23b), we insert the modified element $\hat{a}_{k,1}$ given in (20) and observe that
$\hat{Q}_{k+2}$ $=$ $((a_{k,1}+\alpha)Q_{1}+a_{k,2}Q_{2}+\cdots+a_{k,k+1}Q_{k+1})B_{k+2}$
$=$ $A_{k+2}B_{k+2}+\alpha B_{k+2}=Q_{k+2}+\alpha B_{k+2}.$
To prove (23c), we prove $\hat{A}_{k+3}=A_{k+3},\ldots,\hat{A}_{m+2}=A_{m+2}$, $\hat{B}_{k+3}=B_{k+3},\ldots,\hat{B}_{m+2}=B_{m+2}$, i.e., that the factors for $\hat{Q}_{k+3}$,…, $\hat{Q}_{m+2}$ are unchanged. For $k+3$ we have
(24) $\hat{A}_{k+3}$ $=$ $(\hat{a}_{k+1,1}Q_{1}+\cdots+\hat{a}_{k+1,k+1}Q_{k+1}+a_{k+1,k+2}\hat{Q}_{k+2})$
$=$ $A_{k+3}-\alpha a_{k+1,k+2}(b_{k,1}Q_{1}+\cdots+b_{k,k+1}Q_{k+1})+a_{k+1,k+2}\alpha B_{k+2}$
$=$ $A_{k+3}-\alpha a_{k+1,k+2}B_{k+2}+\alpha a_{k+1,k+2}B_{k+2}=A_{k+3},$
where we have inserted (21a) and (23b) to obtain the second equality. The relation $\hat{B}_{k+3}=B_{k+3}$ can be shown analogously using (21b) and (23b). With induction we can prove the corresponding relation for any factor. Consequently, we have
(25) $\hat{Q}_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{k+2},\quad i=k+1,\ldots,m$
which proves (23c).
The conclusion (22) follows from the same construction as (24).
3.2 Normalized forms and unreduced schemes
The theorems in the previous section can be applied to impose a certain structure on the triplet $(A,B,c)$ without (or with very little) loss of generality. For any triplet $(A,B,c)$ we can invoke Theorem 3.3 repeatedly. By selecting $\alpha$ as the negative element in the first column, we obtain matrices $A$ and $B$ with a first column containing only zeros. The first column corresponds to the addition of a scaled identity, which is independent (constant) with respect to the input $X$. Note that this can be done for any triplet, and no generality is lost by assuming the first column is zero.
Definition 3.5.
If $a_{1,1}=\cdots=a_{m,1}=b_{1,1}=\cdots=b_{m,1}=0$, we call the evaluation scheme a constant-free evaluation scheme.
Corollary 3.6.
Any evaluation scheme is equivalent to a constant-free evaluation scheme.
Similarly, if we assume that the Hessenberg matrices $A$ and $B$ are unreduced [11, pg 381], i.e., the elements $a_{1,2},\ldots,a_{m,m+1}$, $b_{1,2},\ldots,b_{m,m+1}$ are non-zero, we can impose further structure by repeatedly applying Theorem 3.1 with scaling determined by the corresponding last non-zero element of the row. This process imposes a normalization on each row.
Definition 3.7.
A constant-free evaluation scheme satisfying $a_{1,2}=\cdots=a_{m,m+1}=b_{1,2}=\cdots=b_{m,m+1}=1$, is called a normalized evaluation scheme and we call the triplet $(A,B,c)$ normalized.
Corollary 3.8.
Any evaluation scheme corresponding to $(A,B,c)$ where $A$ and $B$ are unreduced Hessenberg matrices, is equivalent to a normalized evaluation scheme.
3.3 Further transformations
For normalized constant-free evaluation schemes, we have $a_{11}=b_{11}=0$ and $a_{12}=b_{12}=1$, meaning that the first multiplication always corresponds to squaring of the input matrix, i.e.,
(26) $Q_{3}=X^{2}.$
This assumption can be made without loss of generality and will be used henceforth. This fact was already observed by Paterson and Stockmeyer [19, pg 61].
We now derive further transformations under a technical assumption. For the first two rows we assume that we have the structure of a unreduced constant-free triplet. [^1] This is assumption is made mostly simplify the derivation, and similar results hold, e.g. when $a_{23}=0$. This in turn is equivalent to assuming $a_{23}=b_{23}=1$.
Consider the following transformation based on modifying the second row. We add $\alpha$ and $\beta$ to the second element in the second row in $A$ and $B$ respectively, i.e.,
(27a) $\hat{Q}_{4}$ $=$ $((a_{22}+\alpha)Q_{2}+Q_{3})((b_{22}+\beta)Q_{2}+Q_{3})$
(27b) $=$ $Q_{4}+(\alpha\beta+\alpha b_{22}+\beta a_{22})X^{2}+(\alpha+\beta)X^{3}.$
If we enforce $\beta=-\alpha$, one term in the modification disappears and we can compensate for the other term in a way such that the $\hat{Q}_{5}=Q_{5},\ldots,\hat{Q}_{m}=Q_{m}$ are unmodified.
Theorem 3.9.
Let $p$ be the polynomial associated with a constant-free triplet $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$, satisfying $a_{23}=b_{23}=1$, and $\hat{p}$ be the polynomial associated with $(\hat{A},\hat{B},\hat{c})$ where
(28a) $\hat{A}$ $=$ $\begin{bmatrix}0&1&&&&&\\ 0&a+\alpha&a&&&&\\ 0&a&\hat{a}_{3,3}&a&&&\\ 0&a&\hat{a}_{4,3}&a&\ddots&&\\ \vdots&\vdots&\vdots&\vdots&\ddots&a&\\ 0&a&\hat{a}_{m,3}&a&\cdots&a&a&\end{bmatrix}$
(28b) $\hat{B}$ $=$ $\begin{bmatrix}0&1&&&&&\\ 0&b-\alpha&b&&&&\\ 0&b&\hat{b}_{3,3}&b&&&\\ 0&b&\hat{b}_{4,3}&b&\ddots&&\\ \vdots&\vdots&\vdots&\vdots&\ddots&b&\\ 0&b&\hat{b}_{m,3}&b&\cdots&b&b&\end{bmatrix}$
and
(29) $\hat{c}=\left[\begin{array}[]{c c c c c c}c_{1}&c_{2}&\hat{c}_{3}&c_{4}&\cdots&c_{m+2}\end{array}\right].$
The modifications to the coefficients are given by
(30a) $\hat{a}_{i,3}$ $=$ $a_{i,3}-sa_{i,4},\quad i=3,\ldots,m.$
(30b) $\hat{b}_{i,3}$ $=$ $b_{i,3}-sb_{i,4}\quad i=3,\ldots,m.$
(30c) $\hat{c}_{3}$ $=$ $c_{3}-sc_{4}.$
where,
$s=-\alpha^{2}+\alpha(b_{22}-a_{22}).$
Then,
(31) $p=\hat{p}.$
Proof 3.10.
From the first row we have $\hat{Q}_{3}=Q_{3}=X^{2}$. The proof is based on establishing the following two statements:
(32a) $\hat{Q}_{4}$ $=$ $Q_{4}+sQ_{3},$
(32b) $\hat{Q}_{i+2}$ $=$ $Q_{i+2},\quad i=3,\ldots,m.$
This, together with the definition of $\hat{c}$ in (30c) implies that the theorem conclusion (31) holds.
To prove (32a), we use the relation (27). More precisely, we substitute $\beta=-\alpha$ into (27b) and obtain
(33) $\begin{split}\hat{Q}_{4}&=Q_{4}+(-\alpha^{2}+\alpha(b_{22}-a_{22}))X^{2}=Q_{4}+sQ_{3}.\end{split}$
To prove (32b), we prove that the constructing factors are unchanged, i.e., $\hat{A}_{3}=A_{3},\ldots,\hat{A}_{m+2}=A_{m+2}$ and $\hat{B}_{3}=B_{3},\ldots,\hat{B}_{m+2}=B_{m+2}$. For the first factor equality we have
$\hat{A}_{5}$ $=$ $a_{32}Q_{2}+\hat{a}_{33}Q_{3}+a_{34}\hat{Q}_{4}$
$=$ $a_{32}Q_{2}+a_{33}Q_{3}-a_{34}sQ_{3}+a_{34}Q_{4}+a_{34}sQ_{3}=A_{5},$
where we have used (30a) and (32a) in the second equality. The relation $\hat{B}_{5}=B_{5}$ can be shown analogously using (30b) and (32a). With induction we can prove the corresponding relation for any factor. Consequently, we have
(34) $\hat{Q}_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{i+2},\quad i=3,\ldots,m,$
which proves (32b). The theorem conclusion (31) follows by the same construction as (3.10).
For a given evaluation scheme satisfying $a_{23}=b_{23}=1$, we can apply the previous theorem with $\alpha=b_{22}$, so that $\hat{b}_{22}=0$. This shows that we can assume $b_{22}=0$ for evaluation schemes that are unreduced in the first two rows, without loss of generality. Under these assumptions we get for the second multiplication:
(35) $Q_{4}=(a_{22}Q_{2}+Q_{3})Q_{3}=a_{22}Q_{2}Q_{3}+Q_{3}^{2}=a_{22}X^{3}+X^{4}.$
Next, we state and prove a theorem for the third row of the coefficient matrices. For this it is necessary to assume that the first three rows are unreduced, i.e., $a_{23}=a_{34}=b_{23}=b_{34}=1$.
Before proceeding, we illustrate the main idea. The theorem is based on perturbing each of the free elements in the third row of $A$ and $B$, while simultaneously preserving the final output polynomial. In particular, we use perturbations with the following structure:
$\hat{a}_{32}$ $=$ $a_{32}+\alpha,$
$\hat{b}_{32}$ $=$ $b_{32}-\alpha,$
$\hat{a}_{33}$ $=$ $a_{33}+\beta,$
$\hat{b}_{33}$ $=$ $b_{33}-\beta.$
It turns out that when the perturbations have this particular structure, $Q_{5}$ is modified by the addition of a linear combination of $Q_{3}$, $Q_{4}$ and $X^{3}$. This is helpful because we can compensate for $Q_{3}$ and $Q_{4}$, since we have access to these matrices directly. We can however not compensate for $X^{3}$ directly. Fortunately, we can ensure that the $X^{3}$-coefficient modification is zero by placing an additional condition on $\alpha$ and $\beta$, encoded in the function $z(\alpha,\beta)=0$.
Theorem 3.11.
Let $p$ be the polynomial associated with the constant-free triplet $(A,B,c)\in\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m\times(m+1)}\times\mathbb{C}^{m+2}$ satisfying $a_{23}=a_{34}=b_{23}=b_{34}=1$, and $\hat{p}$ be the polynomial associated with $(\hat{A},\hat{B},\hat{c})$ where
(37a) $\hat{A}$ $=$ $\begin{bmatrix}0&1&&&&&\\ 0&a&a&&&&\\ 0&a+\alpha&a+\beta&a&&&\\ 0&a&\hat{a}_{4,3}&\hat{a}_{4,4}&\ddots&&\\ \vdots&\vdots&\vdots&\vdots&\ddots&a&\\ 0&a&\hat{a}_{m,3}&\hat{a}_{m,4}&\cdots&a&a&\end{bmatrix}$
(37b) $\hat{B}$ $=$ $\begin{bmatrix}0&1&&&&&\\ 0&0&b&&&&\\ 0&b-\alpha&b-\beta&b&&&\\ 0&b&\hat{b}_{4,3}&\hat{b}_{4,4}&\ddots&&\\ \vdots&\vdots&\vdots&\vdots&\ddots&b&\\ 0&b&\hat{b}_{m,3}&\hat{b}_{m,4}&\cdots&b&b&\end{bmatrix}$
and
(38) $\hat{c}=\left[\begin{array}[]{c c c c c c c}c_{1}&c_{2}&\hat{c}_{3}&\hat{c}_{4}&c_{5}&\cdots&c_{m+2}\end{array}\right].$
The modifications to the coefficients are given by
(39a) $\hat{a}_{i,3}$ $=$ $a_{i,3}-a_{i,5}s_{1},\quad i=4,\ldots,m,$
(39b) $\hat{a}_{i,4}$ $=$ $a_{i,4}-a_{i,5}s_{2},\quad i=4,\ldots,m,$
(39c) $\hat{b}_{i,3}$ $=$ $b_{i,3}-b_{i,5}s_{1},\quad i=4,\ldots,m,$
(39d) $\hat{b}_{i,4}$ $=$ $b_{i,4}-b_{i,5}s_{2},\quad i=4,\ldots,m,$
(39e) $\hat{c}_{3}$ $=$ $c_{3}-c_{5}s_{1}$
(39f) $\hat{c}_{4}$ $=$ $c_{4}-c_{5}s_{2},$
where,
(40a) $s_{1}$ $=$ $\alpha(b_{32}-a_{32})-\alpha^{2}$
(40b) $s_{2}$ $=$ $\beta(b_{33}-a_{33})-\beta^{2}.$
Then,
(41) $p=\hat{p}$
for any $\alpha$ and $\beta$ satisfying
(42) $z(\alpha,\beta)\mathrel{\mathop{:}}=\alpha(b_{33}-a_{33})+\beta(b_{32}-a_{32})-2\alpha\beta-s_{2}a_{22}=0.$
Proof 3.12.
It follows from the first two rows that $\hat{Q}_{3}=Q_{3}=X^{2}$, $\hat{Q}_{4}=Q_{4}=a_{22}X^{3}+X^{4}$. The proof is based on establishing the following two statements:
(43a) $\hat{Q}_{5}$ $=$ $Q_{5}+s_{1}Q_{3}+s_{2}Q_{4},$
(43b) $\hat{Q}_{i+2}$ $=$ $Q_{i+2},\quad i=4,\ldots,m.$
This, together with the definition of $\hat{c}$ in (39e) and (39f) implies that the theorem conclusion (41) holds.
To prove (43a) we will manipulate the expression for $\hat{Q}_{5}$ until we have $Q_{5}$ plus some remainder term. The goal is to show that this remainder term is a linear combination of $Q_{3}$, $Q_{4}$, and $X^{3}$, where the $X^{3}$-coefficient is given by $z(\alpha,\beta)$, which is zero by assumption. If we express $\hat{Q}_{5}$ in terms of the constructing multiplication factors, we get
(44) $\begin{split}\hat{Q}_{5}&=\hat{A}_{5}\hat{B}_{5}\\ &=((a_{32}+\alpha)Q_{2}+(a_{33}+\beta)Q_{3}+Q_{4})((b_{32}-\alpha)Q_{2}+(b_{33}-\beta)Q_{3}+Q_{4})\\ &=\big{(}A_{5}+(\alpha Q_{2}+\beta Q_{3})\big{)}\big{(}B_{5}-(\alpha Q_{2}+\beta Q_{3})\big{)}.\end{split}$
From this we can expand and use that $Q_{5}=A_{5}B_{5}$ to get
(45) $\begin{split}\hat{Q}_{5}&=Q_{5}+(B_{5}-A_{5})(\alpha Q_{2}+\beta Q_{3})-(\alpha Q_{2}+\beta Q_{3})^{2},\end{split}$
which is on the desired form. The next step is to show that the remainder is a linear combination of $Q_{3}$, $Q_{4}$, and $X^{3}$. We factorize the remainder in (45) to obtain
(46) $\begin{split}\hat{Q}_{5}&=Q_{5}+\big{(}B_{5}-A_{5}-(\alpha Q_{2}+\beta Q_{3})\big{)}\big{(}\alpha Q_{2}+\beta Q_{3}\big{)}.\end{split}$
Before proceeding we define
(47) $\begin{split}d_{32}=b_{32}-a_{32},\\ d_{33}=b_{33}-a_{33}.\end{split}$
This allows us to describe the difference between the multiplication factors more compactly:
(48) $\begin{split}B_{5}-A_{5}&=(b_{32}Q_{2}+b_{33}Q_{3}+Q_{4})-(a_{32}Q_{2}+a_{33}Q_{3}+Q_{4})\\ &=(b_{32}-a_{32})Q_{2}+(b_{33}-a_{33})Q_{3}.\\ &=d_{32}Q_{2}+d_{33}Q_{3}.\end{split}$
We substitute this expression into (46) and simplify the first factor of the remainder
(49) $\begin{split}\hat{Q}_{5}&=Q_{5}+\big{(}d_{32}Q_{2}+d_{33}Q_{3}-\alpha Q_{2}-\beta Q_{3}\big{)}\big{(}\alpha Q_{2}+\beta Q_{3}\big{)}\\ &=Q_{5}+\big{(}(d_{32}-\alpha)Q_{2}+(d_{33}-\beta)Q_{3}\big{)}\big{(}\alpha Q_{2}+\beta Q_{3}\big{)}.\\ \end{split}$
Next, we expand the remainder
(50) $\begin{split}\hat{Q}_{5}&=Q_{5}+\alpha(d_{32}-\alpha)Q_{2}^{2}+\big{(}\alpha(d_{33}-\beta)+\beta(d_{32}-\alpha)\big{)}Q_{2}Q_{3}+\beta(d_{33}-\beta)Q_{3}^{2}\\ &=Q_{5}+s_{1}Q_{2}^{2}+(\alpha d_{33}+\beta d_{32}-2\alpha\beta)Q_{2}Q_{3}+s_{2}Q_{3}^{2},\\ \end{split}$
where we have simplified in the last equality using (40a) and (40b). By using $Q_{3}=Q_{2}^{2}$, $Q_{2}Q_{3}=X^{3}$ and $Q_{3}^{2}=X^{4}$, we can rewrite this as
(51) $\hat{Q}_{5}=Q_{5}+s_{1}Q_{3}+(\alpha d_{33}+\beta d_{32}-2\alpha\beta)X^{3}+s_{2}X^{4}.$
Finally, we use (35) in order to express $X^{4}$ in terms of $Q_{4}$ and $X^{3}$
(52) $\begin{split}\hat{Q}_{5}&=Q_{5}+s_{1}Q_{3}+(\alpha d_{33}+\beta d_{32}-2\alpha\beta)X^{3}+s_{2}(Q_{4}-a_{22}X^{3})\\ &=Q_{5}+s_{1}Q_{3}+\big{(}\alpha d_{33}+\beta d_{32}-2\alpha\beta-s_{2}a_{22}\big{)}X^{3}+s_{2}Q_{4}\\ &=Q_{5}+s_{1}Q_{3}+z(\alpha,\beta)X^{3}+s_{2}Q_{4}\\ &=Q_{5}+s_{1}Q_{3}+s_{2}Q_{4},\end{split}$
where we have used that $z(\alpha,\beta)=0$ in the last equality. This proves the statement.
To prove (43b) we show that the constructing factors are unchanged, i.e., $\hat{A}_{6}=A_{6},\ldots\hat{A}_{m+2}=A_{m+2}$ and $\hat{B}_{6}=B_{6},\ldots\hat{B}_{m+2}=B_{m+2}$. For the first factor equality we have
(53) $\begin{split}\hat{A}_{6}&=(a_{42}Q_{2}+\hat{a}_{43}Q_{3}+\hat{a}_{44}Q_{4}+a_{45}\hat{Q}_{5})\\ &=a_{42}Q_{2}+(a_{43}-a_{45}s_{1})Q_{3}+(a_{44}-a_{45}s_{2})Q_{4}+a_{45}Q_{5}+a_{45}(s_{1}Q_{3}+s_{2}Q_{4})\\ &=A_{6}-a_{45}s_{1}Q_{2}-a_{45}s_{2}Q_{4}+a_{45}(s_{1}Q_{3}+s_{2}Q_{4})\\ &=A_{6}.\end{split}$
where we have used (39a), (39b) and (43a) in the second equality. Note that although $a_{45}=1$, writing the coefficient explicitly better demonstrates how to generalize for $\hat{A}_{i+2}=A_{i+2}$, $i>4$, since $a_{i5}\neq 1$ in general. The relation $\hat{B}_{6}=B_{6}$ follows analogously using (39c), (39d) and (43a).
The corresponding relation can be shown for all factors using induction. Consequently, we have
(54) $Q_{i+2}=\hat{A}_{i+2}\hat{B}_{i+2}=A_{i+2}B_{i+2}=Q_{i+2},\quad i=4,\ldots,m.$
which proves (43b). The theorem conclusion (41) follows by the same construction as (53).
Free variables in Theorem 3.11
Note that Theorem 3.11 includes a scalar valued condition $z(\alpha,\beta)=0$, involving two scalar variables. Let $d_{32}=b_{32}-a_{32}$ and $d_{33}=b_{33}-a_{33}$ be defined as in the proof of the theorem. If we let $\alpha$ be given, we get a quadratic equation in $\beta$:
(55) $d_{33}\alpha=a_{22}\beta^{2}+\beta\left(2\alpha+d_{32}-a_{22}d_{33}\right).$
When $a_{22}\neq 0$, the solution $\beta$ to the equation is explicitly availbable from the solution of the quadratic equation. This has a disadvantage of introducing a square root operation. In a real setting, this can yield complex coefficients.
The paper [24] suggest several approaches to evaluate polynomials with a low number multiplications. For the case $m=3$, the formulas [24, equation (31)] involve a square root, and indeed that approach can be derived from the above transformation with $\alpha=-a_{32}$ and solving for $\beta$.
Suppose $\beta$ is given. Then, the solution for $\alpha$ can be expressed as:
(56) $\alpha=\frac{a_{22}\beta^{2}+(d_{32}-a_{22}d_{33})\beta}{d_{33}-2\beta}$
with the condition $\beta\neq\frac{1}{2}d_{33}$. To reframe this condition in terms of entries in matrices $A$ and $B$, we use a change of variables that essentially generalizes [21] where it is given for $m=3$:
(57) $\beta=\frac{1}{2}d_{33}+r,\quad r\neq 0.$
With this choice, we obtain updated table entries:
(58) $\begin{split}\hat{a}_{33}&=a_{33}+\beta=a_{33}+\frac{b_{33}-a_{33}}{2}+r=\frac{a_{33}+b_{33}}{2}+r,\\ \hat{b}_{33}&=b_{33}-\beta=b_{33}-\frac{b_{33}-a_{33}}{2}-r=\frac{a_{33}+b_{33}}{2}-r.\end{split}$
This leads to the relation:
(59) $\hat{a}_{33}-\hat{b}_{33}=2r.$
Consequently, any evaluation scheme that is unreduced in the first three rows can be transformed into an equivalent scheme satisfying:
(60) $b_{33}=a_{33}+1,$
as we can choose any non-zero $r$. Thus, when considering evaluation schemes that are unreduced in the first three rows, we can assume without loss of generality that $b_{33}=a_{33}+1$. This assumption helps reduce the number of variables without introducing additional constraints.
4 Minimality of parameterization
4.1 Minimality of unreduced evaluation schemes
As a consequence of the equivalence theorems in the previous section, we can conclude that any unreduced evaluation scheme can be be assumed to have the form
(61) $[A|B]=\left[\begin{array}[]{ccccccc|cccccccc}0&1&&&&&&0&1&&&&\\ 0&a&1&&&&&0&0&1&&&\\ 0&a&a_{33}&1&&&&0&b&{a_{33}+1}&1&&\\ 0&a&a&a&1&&&0&b&b&b&1&\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&&\vdots&\vdots&\vdots&\vdots&\ddots&\ddots\\ 0&a&a&a&\cdots&a&1&0&b&b&b&\cdots&b&1\\ \end{array}\right],$
with an arbitrary $c$-vector. The dimension of $\Pi_{2^{m}}^{*}$ is determined from the rank of the Jacobian for a generic element in $\Pi_{2^{m}}^{*}$. A generic element in $\Pi_{2^{m}}^{*}$ is unreduced. Hence, we can conclude that the dimension is bounded by the number of free variables. The parameterization (61) contains $m^{2}$ parameters, and therefore, similar to (12), we have the following.
Corollary 4.1.
For $m>2$, we have
$\dim(\Pi^{*}_{2^{m}})\leq m^{2}.$
In order to determine the sharpness of the bound, we have done numerical computations in high precision arithmetic. In our simulations, the numerical rank of the the Jacobian (as defined in Section 2.2) for random $(A,B,c)$ has always coincided with the bound; verified up until $m=10$. Therefore, this is seemingly a minimal parameterization of unreduced schemes in $\Pi_{2^{m}}^{*}$; to our knowledge the first minimal parameterization.
Conjecture 4.2.
For $m>2$, we have
$\dim(\Pi^{*}_{2^{m}})=m^{2}.$
4.2 Reduced evaluation schemes
In practice, the unreduced evaluation schemes are rarely useful to evaluate a given polynomial of a high degree $d$, due to a lack of degrees of freedom in $\Pi_{2^{m}}^{*}$ for large $m$, in comparison to $\dim(\Pi_{2^{m}})=d+1=2^{m}+1$. If the Hessenberg matrices $A$ and $B$ are reduced, we obtain output polynomials of lower degree, for example the pair
(62) $[A|B]=\left[\begin{array}[]{ccccccc|cccccccc}0&1&&&&&&0&1&&&&\\ 0&\times&1&&&&&0&0&1&&&\\ 0&\times&\times&1&&&&0&\times&\times&1&&\\ 0&\times&\times&\times&1&&&0&\times&\times&1&0&\\ 0&\times&\times&\times&\times&1&&0&\times&\times&\times&1&0\\ 0&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&1&0\\ \end{array}\right]$
corresponds to $m=6$ multiplications, but results in a polynomial of degree $32$. With a single reduction, we mean that, in one row we set the last non-zero elements of to zero in either $A$ or $B$. With $r$ reductions we mean the repeated application of a single reduction. Note that we can still normalize each row since the transformation theorems also applies to reduced systems. Hence, we lose a degree of freedom with every reduction, and due to Corollary 4.1, we expect the corresponding dimension to be
(63) $m^{2}-r.$
In the follow section we proceed by studying reduced evaluation schemes, that lead to specific polynomial degrees, and describe ways to compute $(A,B,c)$ for that reduction structure, for a polynomial given in a monomial basis.
[…]
Figure 1: The polynomial degrees when the Hessenberg matrices in the triplet $(A,B,c)$ are reduced Hessenberg matrices. The parameter $\epsilon=0.1$ is selected for visualization reasons. Any point above the blue curve can be discarded to not completely contain the corresponding polynomial subset, in the sense of (64), as it has an insufficient number of degrees of freedom.
5 Polynomial subsets
This section is devoted to the study of the question: What is largest $d$ such that all $d$-degree polynomials can be computed with $m$ multiplications? Formally, we use two versions of this problem
(64) $\max\{d:\Pi_{d}\subset\Pi_{2^{m}}^{*}\}\leq\max\{d:\Pi_{d}\subset\widebar{\Pi_{2^{m}}^{*}}\}.$
We consider the left-hand side when possible, and otherwise study the right-hand side in order to avoid limit cases that can occur similar to (8).
The previous section stressed that we must consider reduced matrices. For a given reduction, we can compute the corresponding polynomial degree. Hence, we can investigate candidate solutions to (64) by considering all possible reductions. This approach is depicted in Figure 1, which illustrates all combinations of reductions for up to $m\leq 7$ multiplications and $r\leq 7$ reductions For instance, (62) corresponds to the scenario (horizontal axis) with $m=6$ multiplications and $r=3$ reductions, and achieves a polynomial degree of $32$, as shown on the vertical axis of the figure. For identifying optimal solutions to (64), higher polynomial degrees are preferable. On the other hand, a too high degree will lead to an insufficient number of degrees of freedom. More precisely, (63) gives a bound on the degree of the admissable candidate solutions to (64) for a given number of multiplications and reductions. This is visualized with a blue line in the figure. Hence, for the purpose of studying (64), we can disregard points above this line.
The problem (64) gets more and more complicated for the higher $m$ is. For $m=3$, the problem is essentially already solved in [24] since an explicit procedure is provided to compute almost all polynomials of degree $8$ with $m=3$ multiplications, i.e.,
$\max\{d:\Pi_{d}\subset\widebar{\Pi_{2^{3}}^{*}}\}=8.$
For $m=4$, similar constructions are also provided in [24], providing a method for degree $12$. Another approach with smaller number of assumptions on the monomial coefficients is provided in Section 5.1. From Figure 1, we see that this is the highest admissible degree and therefore conclude that this is optimal. For $m\geq 5$ we were not able to solve the problem analytically, and instead resort to computational tools. More precisely, we frame the problem with a given reduction and structure as a system of polynomial equations. For $m=5$ we use the package HomotopyContinuation.jl [4], to find seemingly all solutions. For $m=6$ and $m=7$, the system was too complicated, and we were not able to find solutions with this package. However, by using the software [14], we were able to construct locally convergent iterative methods and in all specified test cases, able to find solutions. These simulations combined with reasoning with admissible degrees in Figure 1, we conclude conjectures concerning the solution to (64). For reproducibility, all simulations (including starting values) are provided in the publicly available github repository: https://github.com/GustafLorentzon/polynomial-set-paper.
5.1 Four multiplications
When we study $m=4$, we identify from Figure 1 that the highest degree of the admissible polynomials is $12$, since without reductions we only have $16$ degrees of freedom, which cannot parameterize $\Pi_{16}$. The solution to (64) is indeed 12. This can already be concluded from the method in [24, pg 237] which is a method to evaluate polynomials of degree 12, which are given in their monomial basis, using only $m=4$ multiplications. In our terminology this corresponds to the reduction $a_{2,3}=0$, and additionally $a_{4,2}=0$. The method [24, pg 237] involves square roots of expression containing the monomial coefficients, as well as divisions of certain quantities leading some exceptions corresponding to certain limit cases. Therefore from [24] we conclude that we have
(65) $\max\{d:\Pi_{d}\subset\widebar{\Pi_{2^{4}}^{*}}\}=12$
in a complex sense.
We now present a slightly more general method for evaluating any polynomial in $\Pi_{12}$ in four multiplications, that does not involve square roots or divisions except for the leading monomial coefficient. Consider the evaluation schemes with the following structure
(66f) $\left[\begin{array}[]{c | c}A&B\end{array}\right]$ $=\left[\begin{array}[]{c c c c c | c c c c c}0&1&&&&0&1&&\\ 0&1&0&&&0&0&1&\\ 0&a_{32}&a_{33}&1&&0&0&0&1\\ 0&a_{42}&a_{43}&a_{44}&1&0&b_{42}&b_{43}&a_{44}+1&1\end{array}\right]$
(66h) $c$ $=\left[\begin{array}[]{cccccc}c_{1}&c_{2}&c_{3}&c_{4}&c_{5}\end{array}\right].$
Suppose $\alpha_{0},\ldots,\alpha_{12}$ represent a given polynomial $p(X)=\alpha_{0}I+\alpha_{1}X+\cdots+\alpha_{12}X^{12}\in\Pi_{12}$. If we expand the parameterization we obtain a multivariate polynomial system of equations, one for each monomial coefficient in the output polynomial. In the terminology of Section 2.2 the system corresponds to considering the 00th,$\ldots$,$12$th derivatives of the equation $\Phi(A,B,c)(x)=p(x)$ with respect to $x$ evaluated at $x=0$. In this case we have $13$ equations in the $13$ variables. To solve solve this system, we first introduce the auxiliary variables $\beta_{43}=b_{43}+a_{43}$ and $\beta_{42}=b_{42}+a_{42}$. This system is explicitly solvable by considering the equations in the output polynomial ordered in descending degree, such that the equation corresponding to $\alpha_{12}$ is treated first. The system is in this sense triangular, which was also crucial for the construction in [24, Section 3]. The solution to the system is given by following sequence of equations:
(67a) $c_{6}$ $=\alpha_{12}$
(67b) $a_{33}$ $=\frac{1}{2}\left(\frac{\alpha_{11}}{c_{6}}\right)$
(67c) $a_{32}$ $=\frac{1}{2}\left(\frac{\alpha_{10}}{c_{6}}-a_{33}^{2}\right)$
(67d) $a_{44}$ $=\frac{1}{2}\left(\frac{\alpha_{9}}{c_{6}}-2a_{32}a_{33}-1\right)$
(67e) $\beta_{43}$ $=\frac{\alpha_{8}}{c_{6}}-\left(a_{33}+2a_{33}a_{44}+a_{32}^{2}\right)$
(67f) $\beta_{42}$ $=\frac{\alpha_{7}}{c_{6}}-\left(a_{32}+a_{33}\beta_{43}+2a_{32}a_{44}\right)$
(67g) $c_{5}$ $=\alpha_{6}-c_{6}\left(a_{44}+a_{44}^{2}+a_{33}\beta_{42}+a_{32}\beta_{43}\right)$
(67h) $a_{43}$ $=\frac{c_{5}}{c_{6}}-\left(a_{33}\frac{c_{5}}{c_{6}}+a_{44}\beta_{43}+a_{32}\beta_{42}\right)$
(67i) $a_{42}$ $=\frac{\alpha_{4}}{c_{6}}-\left(a_{32}\frac{c_{5}}{c_{6}}+a_{44}\beta_{42}+a_{43}\beta_{43}-a_{43}^{2}\right)$
(67j) $c_{4}$ $=\alpha_{3}-c_{6}\left(a_{43}\beta_{42}+a_{42}\beta_{43}-2a_{42}a_{43}\right)$
(67k) $c_{3}$ $=\alpha_{2}-c_{6}\left(a_{42}\beta_{42}-a_{42}^{2}\right).$
With the conditions $c_{2}=\alpha_{1}$, $c_{1}=\alpha_{0}$, $b_{43}=\beta_{43}-a_{43}$ and $b_{42}=\beta_{42}-a_{42}$, we have explictly computed all variables in (66).
Recall that $\alpha_{12}\neq 0$ for $p\in\Pi_{12}$, therefore we have made no assumptions other than the degree of the polynomial. Moreover, the formulas preserve the algebraic structure of the variables, e.g., if $\alpha_{0},\ldots,\alpha_{12}\in\mathbb{R}$, then $(A,B,c)$ is a real triplet, i.e., the evaluation coefficients are real. We conclude that
(68) $\max\{d:\Pi_{d}\subset\Pi_{2^{4}}^{*}\}=12$
holds in both a real as well as in a complex sense.
5.2 Five multiplications
For $m=5$ multiplications we see in Figure 1 that the highest degree of the admissible polynomials is $20$. To our knowledge, state-of-the-art is $d=18$ as given in [25, Equation (17)-(19)] with $s=2$ which is based on [24] combined with the Paterson-Stockmeyer evaluation. In our terminology, that approach corresponds to the reduction $a_{23}=a_{56}=0$ and additionally imposing $a_{42}=0$. The reduction in Figure 1 leading to a polynomial of degree $d=20$ corresponds to $a_{45}=a_{56}=0$.
We were not able to explicitly derive a solution to the multivariate polynomial system for the structure $d=20$ analytically, but instead needed to resort to computational tools. The Julia package HomotopyContinuation.jl [4] includes methods to solve polynomial systems of equations based on numerical continuation and with advanced initialization of starting points for the homotopy method. We implemented the system equations stemming from the considering each monomial coefficient for the data structures of this package. Since we have more variables than equations for this structure, we empirically fixed some variables, and chose which variable to fix by trying to reduce the total degree of the system as much as possible. We additionally solved those equations explicitly that could be solved analytically, e.g., first and last equations.
With this setup we were able to compute $(A,B,c)$ for a large number of given polynomials $p\in\Pi_{20}$, including the truncated Taylor expansion of the exponential, as well as the function $1/(1-x)$. The simulation as well as code that computes the corresponding matrix functions are given in the github repository, both for Julia and matlab. For conciseness we report only the numbers for the matrix exponential in the following.
Due numerical condition, we prefer solutions that avoid very large and very small numbers. In our case, to mitigate large numbers we used a scaling of the input. For the exponential we approximate $e^{\alpha x}$ with $\alpha=8$ fixed, since this makes $c_{7}=\alpha^{20}/20!\approx 0.47$ in the order of magnitude one. Note that the scaling of the input can be compensated in the table, although this was not deemed as numerically useful, and hence not used below.
HomotopyConinuation.jl found several solutions and the solution with the smallest values was the following. The parameters were
(69f) $[A|B]$ $=\left[\begin{array}[]{cccccc|cccccc}0&1&&&&&0&1&&&&\\ 0&0&1&&&&0&\frac{1}{2}&1&&&\\ 0&0&2&1&&&0&b&1&1&&\\ 0&a&a&1&0&&0&b&b&b&1&\\ 0&a&a&a&1&0&0&b&b&b&b&1\end{array}\right]$
(69h) $c$ $=\left[\begin{array}[]{cccccccc}c&c&c&c&c&c&c\end{array}\right]$
where the missing values are given in Table 1. The Jacobian of the polynomial system evaluated at this solution has a condition number of $8.1\cdot 10^{2}$.
$a_{42}$ 2.3374451754385963 $c_{1}$ $1$
$a_{43}$ $-41/16$ $c_{2}$ $\alpha$
$a_{52}$ 2.8309861554443847 $c_{3}$6.657689892163032
$a_{53}$ 8.7616118485412 $c_{4}$ 50.445902670306005
$a_{54}$ 5.123957592622475 $c_{5}$ 19.754729172913187
$b_{32}$ 1.4484649122807018 $c_{6}$ 2.8090057997411706
$b_{42}$ 6.389966463262669 $c_{7}$ $\alpha^{20}/20!$
$b_{43}$ 6.697361614351532
$b_{44}$ 2.1451472591988834
$b_{52}$2.458444697550387
$b_{53}$3.6724346694235876
$b_{54}$ 16.044090747953085
$b_{55}$ 7.557067023178642
Table 1: Non-specified values in (69)
Since we were able to find solutions in the case studies we conjecture that this corresponds to a realization of a method for the maximum polynomial degree.
Conjecture 5.1.
$\max\{d:\Pi_{d}\subset\widebar{\Pi_{2^{5}}^{*}}\}=20.$
5.3 Six multiplications
To our knowledge, state-of-the-art for $m=6$ multiplications is $d=24$, again given in [25], with $s=3$. In our terminology, that corresponds to the reduction $a_{23}=a_{33}=a_{34}=a_{67}=0$. Similar to the situation for $m=5$, this is not the highest admissable degree. From Figure 1 we see that the highest admissible degree is $d=32$. With $r=3$, we can use the reduction (62).
Unfortunately, the application of the package HomotopyContinuation.jl was not successful for this case. We have $33$ equations, and $m^{2}-r=33$ unknowns. The creation of initial vectors for the homotopy continuation seems too computationally demanding, possibly due to the high total degree of the polynomial system. Instead we used the package GraphMatFun.jl [14] to create a locally convergent iterative solution method. The system is highly ill-conditioned and a standard Newton approach was not successful. Instead, we employ an iteratively regularized Newton’s method. Following the approach of [6], we compute a Tikhonov-Newton step by applying Newton’s method to a Tikhonov-regularized system. The Tikhonov-Newton step can be computed using the singular value decomposition of the Jacobian matrix, which is directly available from the graph representation in GraphMatFun.jl. The best results were obtained by using Armijo steplength damping and selecting between a Newton and a Tikhonov-Newton step. For the sake of reproducibility, the starting vectors are given explicitly in the software available in the github repository. In order determine conclusively that a solution is found, the solution was post-processed with high precision arithmetic (BigFloat in Julia) such that the 50 first decimals of the coefficients are seemingly accurate.
The above approach with the structure (62) was successful in finding a solution vector for the problem corresponding to the Taylor expansion of the matrix exponential, using complex arithmetic. Unfortunately, we were not able to find a real coefficient vector. The maximum absolute number in the coefficient vector is $3.7\cdot 10^{5}$, and the Jacobian evaluated at the solution is $5.8\cdot 10^{11}$.
From an application viewpoint, real coefficients are advantageous, e.g., since the evaluation of $p(A)$ can be done in real arithmetic if $A$ is real. In order to find a real evaluation scheme, we investigated instead the following structure:
$[A|B]=\left[\begin{array}[]{ccccccc|ccccccc}0&1&&&&&&0&1&&&&\\ 0&\times&1&&&&&0&\times&1&&&\\ 0&\times&\times&1&&&&0&\times&1&0&&\\ 0&\times&\times&\times&1&&&0&\times&\times&\times&1&\\ 0&\times&\times&\times&\times&1&&0&\times&\times&\times&1&0\\ 0&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&1&0\\ \end{array}\right].$
This leads to a polynomial of degree $d=30$. Using the locally convergent iterative method described above, we successfully computed a real solution vector. The maximum absolute number in the coefficient vector is $5.0\cdot 10^{4}$, and the condition number of the Jacobian at the solution is $2.10\cdot 10^{10}$.
The highest degree polynomial set can indeed be different for real and complex case, and the simulations suggest the following.
Conjecture 5.2 (Six multiplications).
For complex coefficients,
$\max\{d:\Pi_{d}\subset{\widebar{\Pi_{2^{6}}^{*}}}\}=32.$
For real coefficients,
$\max\{d:\Pi_{d}\subset{\widebar{\Pi_{2^{6}}^{*}}}\}=30.$
5.4 Seven multiplications
For seven multiplications, we believe state-of-the-art is $d=30$, [25, Table 3] with $s=4$, which in our context corresponds to the reduction $a_{23}=a_{33}=a_{34}=a_{43}=a_{44}=a_{45}=0$ and $a_{78}=0$. In Figure 1, we see that the highest admissible degree seems to be $d=42$ with $r=5$. Using the same technique as for $m=6$ we were able to find a solution to the system for the Taylor expansion of the exponential, if we input scaling $\alpha=16$ and when we used complex arithmetic for the highest admissible degree. We used the following structure:
$[A|B]=\left[\begin{array}[]{cccccccc|ccccccccc}0&1&&&&&&&0&1&&&\\ 0&\times&1&&&&&&0&\times&1&&\\ 0&\times&0&1&&&&&0&\times&1&0&\\ 0&\times&\times&\times&1&&&&0&\times&\times&1&0&\\ 0&\times&0&\times&\times&1&&&0&\times&\times&\times&1&0&\\ 0&\times&\times&\times&\times&\times&1&&0&\times&\times&\times&\times&1&0&\\ 0&\times&\times&\times&\times&\times&\times&1&0&\times&\times&\times&\times&\times&1&0&\\ \end{array}\right].$
The fixed elements in $a_{3,3}=a_{5,3}=0$ were selected empirically to keep the condition number low. The maximum absolute number in the coefficient vector is $2.8\cdot 10^{6}$, and the condition number of the Jacobian at the solution is $2.9\cdot 10^{13}$. From this we conclude the following in a complex sense.
Conjecture 5.3 (Seven multiplications).
$\max\{d:\Pi_{d}\subset{\widebar{\Pi_{2^{7}}^{*}}}\}=42.$
6 Conclusions
This work has a characterization of the set $\Pi_{2^{m}}^{*}$ as its focus, with particular attention given to minimality and computing the maximum degree polynomial subset of $\Pi_{2^{m}}^{*}$ in the sense of (64). The minimality question is from our perspective rather well understood in this paper. The determination of the maximum degree polynomial subset can be further investigated. We only used described $m\leq 7$ with computational reasoning. Both the theoretical description of the general case (e.g. using further tools from algebraic geometry [32]), as well as a more general computational approach could be of interest and useful in practice.
Based on our simulations, one major component is missing before this can be directly used in matrix function evaluation software: Understanding of rounding errors. Evaluation of high-degree polynomials are in general prone to rounding errors, unless special representations such as a Chebyshev basis is used. In this case it seems even more intricate. For example, the coefficients we computed, were computed to full double precision using high precision arithmetic. However, the fact that the system has a rather large condition number, is an indication that in general this evaluation is sensitive with respect to these coefficients. Heuristics similar to [15] might be applicable in a general setting. Further work is needed to determine which evaluation schemes, in the continuuum of $(A,B,c)$ lead to better numerical stability, e.g., in the sense that the condition number is smaller.
The Paterson-Stockmeyer method has proven beneficial not only for evaluating matrix polynomials, but similar computational challenges can arise in other contexts, such as when the input $X$ is a polynomial. In these scenarios, multiplying two quantities is significantly more computationally demanding than forming linear combinations. The construction in the Paterson-Stockmeyer approach resembles the baby-step giant-step (BSGS) technique introduced in [30], which has found various applications and combined with Paterson-Stockmeyer approach in public key and privacy-preserving cryptography [12]. Moreover, both the Paterson-Stockmeyer method and BSGS serve as valuable tools in high-precision arithmetic [16, 31]. Open research questions include how the approach presented in this paper, or methods for $\Pi_{2^{m}}^{*}$ in general, can be applied in these contexts.
The fixed-cost computation approach presented in this paper can be complemented by insights from research on composite polynomials or deep polynomials. See [20] for composite polynomials. Rational approximations corresponding to this concept, such as those in [10], illustrate how successive compositions, e.g., $p(f(g(x)))$, can achieve rapid convergence in terms of both the number of compositions and the parameters involved. Similar findings are noted in [33], motivated by the link between this approach and universal approximation in deep learning. The composite polynomial method can fit within the framework of this paper by zeroing certain elements in matrices $A$ and $B$. Nevertheless, there is a significant distinction in research objectives: our objective is to minimize the number of matrix-matrix multiplications, while [33] focuses on reducing the number of parameters. Although some results, such as the approximation of the $p$th root [10], may be directly applicable, further research is needed to fully explore the differences resulting from these different objectives.
Acknowledgements
The authors are grateful for the discussion and comments from Katlén Kohn, in particular concerning Section 2.2.